OpenStudy (anonymous):

WILL GIVE A MEDAL TO WHOEVER HELPS ME 1. What is the side length b in the triangle below? (1 point)23 16.6 46 32.8

OpenStudy (anonymous):

b^2+20^2=26^2 b^2=676-400 b^2=276 b=sqrt276 (16.61324772583615)

OpenStudy (anonymous):

So a

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can i ask one more maybe 2 more questions

OpenStudy (anonymous):

first send me the triangle

OpenStudy (anonymous):

Just check

OpenStudy (anonymous):

k hold on

OpenStudy (janewelcomes):

Can I ask a question?

OpenStudy (anonymous):

yeah

OpenStudy (janewelcomes):

Awesome. Thank you.

OpenStudy (janewelcomes):

I took a screenshot. I hope this attached

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):